find Approximation transmission probability in quantum mechanics E<U


1
Expert's answer
2021-02-19T18:58:44-0500

Answer

Answer

Let for electron has wave function

"\psi\propto e^\frac{ikx}{\hbar}"

The effective momentum of electron is given by


"\frac{\hbar^2 k^2}{2m*}=-q\phi"

"k=i\sqrt{\frac{2m*q\phi}{\hbar^2}}"

So wavefunction become

"\psi\propto e^{i\sqrt{\frac{2m*q\phi}{\hbar^2}}x}"


"\frac{\psi (x=d) }{\psi(x=0) }\propto \frac{ e^{-\sqrt{\frac{2m*q\phi}{\hbar^2}}.d}}{e^{-\sqrt{\frac{2m*q\phi}{\hbar^2}}.0}}\propto

e^{-\sqrt{\frac{2m*q\phi}{\hbar^2}}d}"


Let JI

be the incident current and JT

be the transmitted current. Then


Transmission probability


"T\propto \frac{J_T}{J_0}\propto (\frac{\psi(x=d) }{\psi(x=0) }) ^2\\\propto e^{-2\sqrt{-\frac{2m*q\phi}{\hbar^2}}d}"